OpenStudy (anonymous):

Point G is (2,9) on the graph of f(x) = 3^x, as shown. The coordinates of point G under the transformation f(-2x), are

OpenStudy (anonymous):

looked this up. for f(nx) transformations you scale by 1/n. for f(-x) transformations you flip across y-axis. ...still haven't confirmed proper way to combine was the earlier answer wrong?

OpenStudy (anonymous):

i think so, i just need 100% help because i need to get a 100 on this assignment

OpenStudy (anonymous):

gotcha. I know the function transformation is correct, but I can't recall exactly where a previous point like G would fall on the new function

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I mean, point G is obviously the second range value (x = 2) .....which is why i considered -2, 71...

OpenStudy (anonymous):

right but i dont understand why y is so high of a number

OpenStudy (anonymous):

because the transformation is doubling the exponent

OpenStudy (anonymous):

o

OpenStudy (anonymous):

idk u might be right

OpenStudy (anonymous):

3^2 = 9 right? well 3^(-2*-2) = 3^4 = 81**** btw

OpenStudy (anonymous):

ooooo

OpenStudy (anonymous):

anyway good luck- hope you get it!

OpenStudy (anonymous):

ok